OpenStudy (jennamini):

Express the given function h as a composition of two functions f and g so that h(x)=(f o g)(x), where one of the functions is x^3-6. H(x)= ^7 square root of x^3-6

OpenStudy (sweetburger):

is that the 7th root of x^3-6?

OpenStudy (jennamini):

the last part looks like this. \[h(x)= \sqrt[7]{x^3-6}\]

OpenStudy (jennamini):

yes!

OpenStudy (sweetburger):

I think you could write it like this \[f(x)=\sqrt[7]{x} \] and \[g(x) = x^3-6\]

OpenStudy (sweetburger):

f(g(x))= I think h(x)

OpenStudy (jennamini):

oh okay so to get the answers for each you kind of just separate the equation?

OpenStudy (sweetburger):

Pretty much. They tell you one equation has to be x^3-6 so you just have to figure out how to transfer that to 7th root (x^3-6)

OpenStudy (jennamini):

oh ok... thank you!!!! :)

OpenStudy (sweetburger):

np glad to help